A block of mass 8.37 kg is released from rest from point A and slides on the track shown. Assume that as the mass slides fromAnd determine the speed of the block at B.
Answer #1

cising AU+AK= WG 4K= Wf-AU W = 7585 AU = 8-37%9*( 3.200-5oo)m = -147.85 AK = -75-8+147-8J = 72J * If the speed at a I zen, Ka